# A Survey on Latent Reasoning

### v1

@@ 3.2 Horizontal Recurrent: Hidden state-based Methods
Traditional approaches like Transformers handle temporal information by storing previous token inputs as key-value pairs in the KV cache. While this preserves most information, it suffers from unbounded memory consumption that scales linearly with sequence length. To address this challenge, we can compress previous information into a fixed-size vector or matrix, similar to RNNs. When working with hidden states, there are two primary approaches to enhance their expressiveness: (1) the Linear-State recurrence approach, which applies update and decay rules to the hidden states, and (2) Gradient-State recurrence approach, treating hidden states as online-learning parameters and optimizing them using online learning methods. Notably, although these methods have not yet produced evidence demonstrating enhanced reasoning capabilities, their theoretical properties suggest they may play a significant role in the future, as they represent a form of iterative processing that is conceptually similar to layer stacking.
@@ 3.2.2 Gradient-State Recurrence
This insight paved the way for a second research trajectory that abandons closed-form descriptions entirely, in favor of direct online learning formulations [5, 3, 97, 52, 4]. This line of work, progressing from TTT (implementing SGD-like dynamics) [97] to Titans (incorporating Adam-like behaviors) [3] and ATLAS (utilizing Muon optimization principles) [4], formulates the state update explicitly as a gradient-based optimization step. Despite their different origins, these approaches converge conceptually and can be understood through the general update rule:
@@ 4.2 Mechanisms of Latent CoT in Layer Representation
Intermediate layers have a causal influence on final reasoning outcomes. Correct activation of these layers is necessary for the model to produce valid inferences. A series of studies identify specialized neurons in intermediate layers and perform causal interventions. They find that enhancing activations significantly improves reasoning performance, while suppressing activations leads to a decline in reasoning ability [108, 31]. Intermediate layer representations, acting as bridging entities, also play a causally critical role in multi-step reasoning outcomes [115]. The functional specialization of intermediate layers makes their correct activation critically decisive for the final reasoning outcomes.
@@
Given the layer specialization, the flow of information across these layers is crucial for reasoning process. Stolfo et al. [94] quantify the indirect contributions of MLP and attention modules to clarify internal information flow pathways in LLM during arithmetic tasks. The results highlight the crucial role of the attention mechanism in inter-layer information flow during reasoning, which transmits computational information from early processing layers to the final token. Wang et al. [103] discover a “generalizing circuit” emerging during the grokking process. This circuit enables cross-layer information flow, with lower layers extracting bridge entities and higher layers conducting reasoning. Yu et al. [124] present a neuron-level investigation into the logits flow of LLMs during multi-hop knowledge prediction. With "back attention" mechanism, hidden information can be effectively transmitted from higher layers to lower layers, enhancing model’s reasoning ability.
@@ 5.2.1 Towards an ’Infinitely Long’ Optimizer Network
Munkhdalai et al. [70] attach a compressive memory to every Transformer block. Each incoming segment updates this memory via a linear–delta rule that asymptotically approaches the fixed-point of an associative array, allowing the model to stream infinitely long inputs with $\mathcal{O}(1)$ memory. From the optimization viewpoint, the Infini-memory applies an online regression step to match past keys to their values, while its gated aggregation with local causal attention preserves short-range precision.

### v2

@@ 3.2 Horizontal Recurrent: Hidden state-based Methods
The standard Transformer provides a baseline for this horizontal expansion. It handles temporal information by storing all previous token inputs as key-value pairs in what is known as the KV cache. This cache effectively serves as the model’s hidden state, preserving a rich history of the sequence. However, this approach has a critical bottleneck: the KV cache grows linearly with the sequence length, leading to unbounded memory consumption that makes processing very long sequences impractical.
To address this challenge, we can compress previous information into a fixed-size vector or matrix, similar to RNNs. When working with hidden states, there are two primary approaches to enhance their expressiveness: (1) the Linear-State recurrence approach, which applies update and decay rules to the hidden states, and (2) Gradient-State recurrence approach, treating hidden states as online-learning parameters and optimizing them using online learning methods. Notably, although these methods have not yet produced evidence demonstrating enhanced reasoning capabilities, their theoretical properties suggest they may play a significant role in the future, as they represent a form of iterative processing that is conceptually similar to layer stacking.
@@ 3.2.2 Gradient-State Recurrence
This insight paved the way for a second research trajectory that abandons closed-form descriptions entirely, in favor of direct online learning formulations [5, 3, 100, 52, 4]. This line of work, progressing from TTT (implementing SGD-like dynamics) [100] to Titans (incorporating Adam-like behaviors) [3] and ATLAS (utilizing Muon optimization principles) [4], formulates the state update explicitly as a gradient-based optimization step. Extending this optimization perspective, Ref. [58] introduce LATENTSEEK, a framework that performs test-time instance-level adaptation by directly optimizing latent representations using policy gradient.Despite their different origins, these approaches converge conceptually and can be understood through the general update rule:
@@ 4.2 Mechanisms of Latent CoT in Layer Representation
Intermediate layers have a causal influence on final reasoning outcomes. Correct activation of these layers is necessary for the model to produce valid inferences. A series of studies identify specialized neurons in intermediate layers and perform causal interventions. They find that enhancing activations significantly improves reasoning performance, while suppressing activations leads to a decline in reasoning ability [111, 31]. Intermediate layer representations, acting as bridging entities, also play a causally critical role in multi-step reasoning outcomes [120]. The functional specialization of intermediate layers makes their correct activation critically decisive for the final reasoning outcomes. For example, Ref. [62] traced failures in multi-hop reasoning to specific Attention modules in the intermediate layers that improperly handled implicit reasoning steps. By successfully "patching" these modules to correct the reasoning, they provided strong causal evidence for the functional specialization of these intermediate layer circuits.
@@
Given the layer specialization, the flow of information across these layers is crucial for reasoning process. Stolfo et al. [97] quantify the indirect contributions of MLP and attention modules to clarify internal information flow pathways in LLM during arithmetic tasks. The results highlight the crucial role of the attention mechanism in inter-layer information flow during reasoning, which transmits computational information from early processing layers to the final token. Wang et al. [106] discover a “generalizing circuit” emerging during the grokking process. This circuit enables cross-layer information flow, with lower layers extracting bridge entities and higher layers conducting reasoning. Yu et al. [129] present a neuron-level investigation into the logits flow of LLMs during multi-hop knowledge prediction. With "back attention" mechanism, hidden information can be effectively transmitted from higher layers to lower layers, enhancing model’s reasoning ability. Further research substantiates this by analyzing the "embedding trajectory" across all model layers. One study [113], which terms this the "Chain-of-Embedding," shows that the trajectory’s geometric shape can distinguish correct from incorrect answers, enabling output-free self-evaluation. Another study [112] uses trajectory "volatility" to detect out-of-distribution mathematical problems, finding that models show an "Early Stabilization" in their reasoning path for familiar tasks but not for unfamiliar ones. Both studies confirm that the vertical, layer-by-layer processing of LLMs contains a rich, interpretable information flow analogous to a latent chain of thought.
@@ 5.2.1 Towards an ’Infinitely Long’ Optimizer Network
Munkhdalai et al. [72] attach a compressive memory to every Transformer block. Each incoming segment updates this memory via a linear–delta rule that asymptotically approaches the fixed-point of an associative array, allowing the model to stream infinitely long inputs with $\mathcal{O}(1)$ memory. However, a reproduction [73] attempt documented significant practical challenges with this approach. Their key finding was that the model’s long-context performance degraded as the number of memory compression steps increased. The authors also reported severe convergence issues, particularly with the gating parameters that balance local and compressed memory, ultimately concluding that the technique was not reliable for extending pretrained models. This empirical evidence suggests that while the idea of memory compression is promising, the specific mechanism in Infini-attention may not be effective in practice, and other methods like rope scaling or Ring Attention are currently more viable options.
@@
In contrast to methods relying on frequent, small-batch updates, recent work [136] argues that this strategy suffers from severe computational inefficiency due to low hardware utilization. The proposed solution, Large Chunk Test-Time Training (LaCT), advocates for the opposite: updating "fast weights" using extremely large chunks of data, ranging from thousands to over a million tokens. This large-chunk paradigm dramatically improves hardware utilization without custom kernels and, more importantly, enables the scaling of nonlinear state sizes to a much larger fraction of the model’s parameters (up to 40%). This enhanced state capacity, combined with sophisticated optimizers like Muon, has been validated across diverse tasks, including novel view synthesis, language modeling, and autoregressive video diffusion.